OpenStudy (anonymous):

Find a polynomial function with real coefficients that has the given zeros. (There are many correct answers.) 2, −2i

zepdrix (zepdrix):

oh oh oh thinking

zepdrix (zepdrix):

I'm not sure of how to do this systematically.. I was just trying to think of something.. then it said it had to be a polynomial.. hmm But finally, :O i think this works, lemme know what you think. f(x) = (x-2)(x^2 + 4)

OpenStudy (unklerhaukus):

\[(x+ 2i)(x-2i)=0\]\[x^2+4=0\]

OpenStudy (anonymous):

I attempted that, webassign didn't like it. I got the question wrong, but my test is tuesday and I'd really like to figure this out.

OpenStudy (anonymous):

unkle how does that work?! where would you then get the positive two from, I can see the -2i...

zepdrix (zepdrix):

that's what im wondering :D heh

OpenStudy (unklerhaukus):

\[f(x)=x^2+4=0\] \[f(-2i)=(-2i)^2+4=-4+4=0\]\[f(2i)=(2i)^2+4=-4+4=0\] oh i thought the roots were \(2i,-2i\)

zepdrix (zepdrix):

Oh double, try this maybe? :o \[x^3 - 8 = 0\]

OpenStudy (anonymous):

but its 2 and -2i, only one has an imaginary number.

zepdrix (zepdrix):

x^3 - 8 should have 3 roots, 1 real and 2 imaginary 2, 2i, -2i I think at least :o

OpenStudy (anonymous):

Zep's first polynomial works just fine. (Second one doesn't though. (2i)^3 - 8 = -8i - 8. You aren't going to be able to have anything smaller than a cubic. Complex roots come in pairs (the root and its conjugate).

zepdrix (zepdrix):

Hmm darn :d

OpenStudy (unklerhaukus):

\[f(x)=(x-2)(x+2i)=x^2-2x+2ix-4i\] \[f(2)=((2)-2)((2)+2i)=0(2+2i)=0\]\[f(-2i)=((-2i)-2)((-2i)+2i)=(-2i-2)0=0\]

zepdrix (zepdrix):

That polynomial doesn't have real coefficients Uncle :'O

OpenStudy (anonymous):

unkle that's what I put as my first answer as well. it was wrong. it seemed logical to me. simple roots. and zep, I used the first one as well, that wasn't accept, the second IDK

OpenStudy (tamtoan):

use original zepdrix answer : (x -2)(x^2 +4) = 0 x^3 - 2x^2 + 4x -8 = 0 <===

OpenStudy (anonymous):

tam, that was the answer that killed me. answers I got, x^3 -2x^2-4x+8 ... x^3-2x^2+2x-4 ... I tried putting them as the zeros themselves ....

OpenStudy (anonymous):

either way I'm out of guesses on the hw, and I cannot seem to figure this out to save my life, so I guess I'll accept this one wrong on the test tuesday.

OpenStudy (tamtoan):

x = -2i ==> x^2 = 4i^2 = -4 ==> x^2 +4 = 0 x = 2 ==> x -2 = 0 so P(x) = 0 = (x - 2)(x +2i) = (x -2)(x^2 +4) = x^3 - 2x^2 + 4x - 8 i don't know where the logic fail so that the answer wasn't right, check with your teacher again ? if you got other explanation from your teach, let us all hear and learn from it. thanks

OpenStudy (tamtoan):

your teacher i mean.